### v2

# Better Fair than Sorry: Adversarial Missing Sensitive Value Imputation for Fair GNNs

Abstract
This paper addresses the problem of learning fair Graph Neural Networks (GNNs) under missing protected attributes. GNNs have achieved state-of-the-art results in many relevant tasks where decisions might disproportionately impact specific communities. However, existing work on fair GNNs assumes that either protected attributes are fully observed or that the missing protected attribute imputation is fair. In practice, biases in the imputation will be propagated to the model outcomes, leading them to overestimate the fairness of their predictions. We address this challenge by proposing Better Fair than Sorry (BFtS), a fair missing data imputation model for protected attributes used by fair GNNs. The key design principle behind BFtS is that imputations should approximate the worst-case scenario for fairness—i.e. when optimizing fairness is the hardest. We implement this idea using a 3-player adversarial scheme where two adversaries collaborate against a GNN-based classifier to maximize its bias. Experiments using synthetic and real datasets show that BFtS often achieves a better fairness $\times$ accuracy trade-off than existing alternatives.
1 Introduction
With the increasing popularity of machine learning models for high-stakes decision-making, it has become a consensus that (1) these models carry implicit biases and (2) these biases should be addressed to improve the fairness of algorithmic decisions Ghallab (2019). The disparate treatment of such models towards African Americans and women has been illustrated in the well-documented COMPAS Angwin et al. (2022) and Apple credit card Vigdor (2019) cases. While there has been extensive research on fair algorithms and fair machine learning, the proposed solutions have mostly disregarded important challenges that arise in real-world settings.
(a) Missing values.
(b) Imputed sensitive attributes.
(c) Worst-case imputation.
Figure 1: Motivation: In Figure (a), a graph machine learning algorithm is applied to decide who receives credit (positive or negative) based on a possibly missing sensitive attribute, gender (and binary only for illustrative purposes). As shown in Figure (b), traditional missing data imputation does not account for outcomes (positive/negative) and thus their imputed values can under-represent the bias of the complete dataset—demographic parity (DP) is $0.09$ in this example. This paper proposes BFtS, an imputation method for graph data that optimizes fairness in the worst-case imputation scenario using adversarial learning, as shown in Figure (c), where DP is $0.47$.
Existing work in fair machine learning has focused on tabular, image, and text data Mehrabi et al. (2021); Pessach and Shmueli (2022). However, in several applications, data can be naturally modeled as graphs (or networks), representing different objects, their relationships, and attributes Barabási (2016); Newman (2018). Graph Neural Networks (GNNs) have achieved state-of-the-art results in many graph machine learning tasks, including node classification, link prediction, and graph classification Gilmer et al. (2017); Kipf and Welling (2017); Veličković et al. (2017a); Hamilton et al. (2017). For instance, in the case of link prediction in a professional network, such as LinkedIn, which is a key recruiting and networking tool, link recommendations should not be biased against protected groups Fabbri et al. (2020); Karimi et al. (2018); Current et al. (2022). However, guaranteeing fairness in graph data is a challenge due to well-known correlations in the network caused by homophily and influence. Anagnostopoulos et al. (2008); Aral et al. (2009).
An additional challenge for fair machine learning is the pervasiveness of missing data in many applications Little and Rubin (2019); Buck (1960); Enders (2010). This is particularly critical because fair algorithms often require knowledge of protected (or sensitive) attributes that are more likely to be missing due to biases in the collection process or privacy concerns (e.g., gender, race, age). As an example, data absence in census surveys is correlated with sensitive attributes such as race, etc. O’Hare and O’Hare (2019); Warren and Passel (1987). Similarly, gender information is often missing in health surveys Weber et al. (2021); Molenberghs and Kenward (2007). Missing data is also a concern in networked data, especially when tracking the transmission of diseases Huisman (2009); Gile and Handcock (2006, 2017); Ghani et al. (1998). As a consequence, many fair machine learning algorithms require the use of missing imputation methods to mitigate their biases, and biases in imputed values might be inadvertently propagated to the final decisions Mansoor et al. (2022).
This paper addresses the problem of mitigating bias in graph machine-learning algorithms in the presence of missing sensitive attributes. We propose an adversarial imputation method that generates challenging instances for a broad class of fair algorithms based on GNNs. Figure 1 motivates our approach. We use a machine learning algorithm that decides whether individuals should or should not receive credit. Figure 0(a) shows both the gender and outcome of each individual. The genders of some samples are unknown in Figure 0(a) and the demographic parity (DP) considering only the observed values is $0.25$ (DP is a bias metric described in Section 3.3). We illustrate how different missing data imputations affect the fairness of credit decisions in Figures 0(b) and 0(c). A straightforward imputation is demonstrated in Figure 0(b), where the gender of the majority of neighbors is assigned to the missing attribute. This results in a best-case scenario in terms of fairness with a DP of $0.09$. However, in the worst-case scenario, shown in Figure 0(c), DP is $0.47$. The bias of the dataset might be highly underrepresented in the imputation of Figure 0(b). The final decision of a fair algorithm will still be unfair if the imputation underestimates the bias. Our empirical results provide strong evidence that off-the-shelf imputation algorithms often misguide fair graph machine learning algorithms due to underestimated values of bias.
We propose Better Fair than Sorry (BFtS), which is a 3-player adversarial learning framework for missing data imputation that produces worst-case imputations for a fair GNN. Figure 2 provides an overview of the proposed approach. The GNN $f_{C}$ interacts with two adversaries, $f_{A}$ and $f_{I}$. $f_{A}$ attempts to predict sensitive values based on the representations learned by $f_{C}$ while $f_{I}$ imputes missing values that help $f_{A}$ maximize its accuracy. As high accuracy of $f_{A}$ indicates high bias in the representations learned by $f_{C}$, $f_{C}$’s parameters are optimized to minimize the bias of its representations in the worst-case in terms of fairness. Our experiments show that BFtS achieves a better accuracy $\times$ fairness trade-off than existing approaches using real datasets and synthetic datasets.
We summarize the contributions of this paper as follows:
  • We investigate a novel problem: fair imputation of missing sensitive data for graph machine learning.
  • We propose Better Fair than Sorry (BFtS), a novel 3-player adversarial learning framework for the imputation of missing sensitive data that produces challenging values for graph-based fairness.
  • We conduct a theoretical analysis showing that BFtS is more robust to convergence issues in training.
  • We demonstrate that BFtS achieves a better fairness $\times$ accuracy trade-off than alternative solutions.
1.1 Related Work
In this section, we will overview related work regarding missing data imputation and fairness in graph machine learning.
1.1.1 Fairness in graph machine learning
We divide the work in this area into three main directions (pre-processing, extended objective functions, and adversarial learning). FairOT and FairDrop are pre-processing techniques that remove the bias in the graph or the node attributes before training the GNN Laclau et al. (2021); Spinelli et al. (2021). Fairwalk, Crosswalk, Debayes MONET, NIFTY, FairVGNN, and FairGAE modify the objective function of GNNs to learn fair representations during training Rahman et al. (2019); Khajehnejad et al. (2022); Buyl and De Bie (2020); Palowitch and Perozzi (2019); Agarwal et al. (2021); Wang et al. (2022); Fan et al. (2021). CFC, FLIP, DKGE, and Debias train adversarial networks along with GNN for fair prediction Bose and Hamilton (2019); Masrour et al. (2020); Arduini et al. (2020); Zhang et al. (2018). However, notice that different from our work, all these methods require fully observed sensitive attributes to ensure fairness.
1.1.2 Missing Data Imputation and fairness
Missing data can be imputed using unconditional mean, reconstruction, preferential attachment, autoencoders, etc. Donders et al. (2006); Pereira et al. (2020); Huisman (2009). The traditional procedure for handling missing data is to first impute the missing values, and then solve the task Rossi et al. (2021); Buck (1960). SAT is a matching-based GNN for graphs with missing attributes Chen et al. (2020) but it does not account for fairness. Training a classifier from imputed data can amplify the unfairness of a machine learning model, as discussed in Subramonian et al. (2022); Zhang and Long (2021); Feng et al. (2023). Some studies try to generate fair feature imputations. Subramonian et al. (2022) has modified mean aggregation feature imputation to minimize discrimination risk against protected groups. Feng et al. (2023) has proposed recursive partitioning and fair uniform resampling for fair classification with missing values. However, as the approaches discussed earlier, these methods require full knowledge of the protected attributes.
1.1.3 Fairness with Missing Protected attributes
These studies are the ones most related to our work. One extreme assumption is the complete unavailability of sensitive information. For instance, Hashimoto et al. (2018) takes into account the worst-case distribution across a given group size, and Lahoti et al. (2020) adversarially reweighs training samples to optimize the average loss. Chai and Wang (2022) employs top $k$ average loss minimization to guarantee fairness. However, in practice, some sensitive information is often available and relatively little information can also significantly improve fairness Chai and Wang (2022). A more practical approach is, therefore, assuming that limited sensitive information is available. FairGNN Dai and Wang (2021), which is the most related work to our work, learns fair Graph Neural Network (GNN) representations while handling missing sensitive attributes in the context of node classification. The key differences between our approach (BFtS) and FairGNN are the following: (1) we propose a fair missing protected value imputation method using a 3-player scheme to generate worst-fairness imputations; (2) we propose using the LDAM loss Cao et al. (2019) instead of the cross-entropy loss to handle imbalances in the data; and (3) we consider a more comprehensive set of datasets in our evaluation. Nguyen et al. (2017); Vandenhende et al. (2019) have applied 3-player adversarial networks to improve classification and the stability of adversarial learning. We also compare our method against RNF Du et al. (2021), which is a more recent approach that generates proxy annotations for missing sensitive attributes by training the classification network with generalized cross-entropy loss Zhang and Sabuncu (2018). Our experiments, based on four real and one simulation datasets, show that BFtS often outperforms FairGNN and RNF in terms of the fairness $\times$ accuracy trade-off.
2 Fairness-Aware Adversarial Missing Data Imputation for Graphs
In this section, we introduce a graph learning framework, which we call BFtS (Better Fair than Sorry), that allows fair GNN training in a settings with missing sensitive data. BFtS applies a 3-player adversarial learning scheme, where the fair GNN is trained jointly with two adversaries—one to predict sensitive attributes based on GNN embeddings and another to impute missing values that minimize fairness. As a result, fairness assessments by the GNN are based on the worst-case imputed values generated by the adversarial imputation.
2.1 Notation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$ be an undirected graph where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, and $\mathcal{X}$ are node attributes. The matrix $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of $\mathcal{G}$ where $A_{uv}=1$ if there is an edge between nodes $u$ and $v$ and $A_{uv}=0$, otherwise. Let the sensitive (or sensitive) attribute be $s$. We assume the semi-supervised scenario, where $\mathcal{V}_{L}\subseteq\mathcal{V}$ nodes are labeled, with binary labels $y_{v}\in\{0,1\}$. We focus on the setting where sensitive attributes might be missing and only $\mathcal{V}_{S}\subseteq\mathcal{V}_{L}$ nodes include the information of the sensitive attribute $s_{v}$ for each node $v$. To simplify the discussion, we will assume that $s_{v}\in\{0,1\}$. Finally, we will consider the node classification task, where each node $v$ has a class label $y_{v}$. We notice that all these assumptions have as their main purpose to facilitate exposition and can be generalized.
2.2 Graph Neural Network
Two of the models (or players) learned by BFtS are GNNs. Without loss of generality, we will assume that they are message-passing neural networks Gilmer et al. (2017), which we will simply refer to as GNNs. A GNN learns node representations $\mathbf{h}_{v}$ for each node $v$ that can be used to predict (categorical or continuous) values $y_{v}=f(\mathbf{h}_{v})$. These representations are learned via successive $\mathsf{AGGREGATE}$ and $\mathsf{COMBINE}$ operations defined as follows:
$\mathbf{h}_{v}^{k}\!=\!\mathsf{COMBINE}^{k}(\mathbf{h}_{v}^{k-1},\mathsf{AGGREGATE}^{k-1}({\mathbf{h}_{u}^{k-1},A_{u,v}\!=\!1}))$
where $\mathbf{a}_{v}^{k}$ aggregates the messages received by the neighbors of $v$ in $\mathcal{G}$ and $\mathbf{h}_{v}^{k}$ are representations learned at the $k$-th layer.
Figure 2: Proposed 3-player framework for fair GNN training with missing data imputation (BFtS). $f_{C}$ generates node representations by minimizing the classification loss $\mathcal{L}_{C}$ (Eqn. 1) and maximizing sensitive attribute prediction loss $\mathcal{L}_{A}$ (Eqn. 3). $f_{A}$ predicts sensitive attributes using the representations from $f_{C}$ by minimizing $\mathcal{L}_{A}$. $f_{I}$ predicts missing values by minimizing the imputation loss $\mathcal{L}_{I}$ (Eqn. 2) and maximizing $\mathcal{L}_{A}$. $\hat{y}$, $\hat{si}$, $\hat{sa}$ are the predictions of $f_{C}$, $f_{I}$ and $f_{A}$, respectively.
2.3 Proposed Model (BFtS)
Our goal is to learn a fair and accurate GNN $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ for node classification with missing sensitive values. The proposed solution is based on adversarial learning and formulates the imputation model as an second adversary of the GNN.
Figure 2 depicts the flow diagram of the proposed model. The model has three primary components: a missing sensitive attribute imputation GNN $f_{I}$, a node classification GNN $f_{C}$, and a sensitive attribute Deep Neural Net(DNN) prediction $f_{A}$. $f_{C}$ takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and predicts the node labels. $f_{A}$ is an adversarial neural network that attempts to estimate the sensitive information from the final layer representations of $f_{C}$ to assess the representations’ bias. More specifically, $f_{C}$ is assumed to be biased if the adversary $f_{A}$ can accurately predict the sensitive attributes information from the representations of $f_{C}$. To train $f_{C}$, we need to know the values of the missing sensitive attributes. $f_{I}$ predicts the missing sensitive attributes. It takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and generates the missing sensitive attributes $\hat{si}$. $f_{I}$ aims to generate sensitive values that minimize the fairness of $f_{C}$. Therefore, it works as a second adversary to the model.
We will first formalize the architecture of the players, then define the their loss functions, and finally describe the training procedure for our approach.
2.3.1 Player architectures
We describe the architecture of each player in BFtS.
GNN classifier $f_{C}$:
This is a node classification model $\hat{y}_{v}=f_{C}(\mathcal{X}_{v},\mathcal{G})$ implemented using a fair GNN. We assume that $f_{C}$ does not apply sensitive attributes for its training but uses other attributes in $\mathcal{X}$ that might be correlated with sensitive ones. The goal is for $f_{C}$ to be both accurate and fair. To improve fairness, $f_{C}$ tries to minimize the loss of adversary $f_{A}$.
sensitive attribute predictor $f_{A}$:
This is a neural network that applies latent node representations $\mathbf{h}_{v}$ generated by $f_{C}$ to predict sensitive attributes as $\hat{sa}_{v}=f_{A}(\mathbf{h}_{v})$. The idea is that $f_{C}$ is fair if $f_{A}$ performs poorly.
Missing data imputation GNN $f_{I}$:
It predicts missing sensitive attributes as $\hat{si}_{v}=f_{I}(\mathcal{X}_{v},\mathcal{G})$. However, besides being accurate, $f_{I}$ also plays the role of an adversary to $f_{C}$ by predicting values that maximize the accuracy of $f_{A}$.
2.3.2 Loss Functions
We describe the loss functions applied in our approach. As most real datasets in the fairness literature suffer from data imbalance, we use the Label-Distribution-Aware Margin (LDAM) loss Cao et al. (2019) for classification and imputation.
Node Classification Loss
Let $\hat{y}=f_{C}(X,G)$ be the one hot label predictions and $\hat{y}^{j}_{v}$ be the jth output of node $v$ for the jth class. Then, the LDAM loss of $f_{C}$ is defined as:
$\mathcal{L}_{C}=\frac{1}{|\mathcal{V_{L}}|}\sum_{v=1}^{|\mathcal{V}_{L}|}-log\frac{e^{\hat{y}^{j}_{v}-\Delta_{j}}}{e^{\hat{y}^{j}_{v}-\Delta_{j}+\sum_{k\neq j}e^{\hat{y}^{k}_{v}}}},$ (1)
where $\Delta_{j}=C/n_{j}^{\frac{1}{4}}$. Here $C$ is a hyperparameter and $n_{j}$ is the number of training samples that belong to class $j$.
Sensitive Attribute Imputation Loss
We also use the LDAM loss to learn the parameters of $f_{I}$:
$\mathcal{L}_{I}=\frac{1}{|\mathcal{V_{S}}|}\sum_{v=1}^{|\mathcal{V}_{S}|}-log\frac{e^{\hat{si}^{j}_{v}-\Delta_{j}}}{e^{\hat{si}^{j}_{v}-\Delta_{j}+\sum_{k\neq j}e^{\hat{y}^{k}_{v}}}}$ (2)
Sensitive Attribute Prediction Loss
We have the sensitive information of $\mathcal{V}_{s}$. We first replace the $\hat{si}_{v}$ with the $\hat{s}_{v}$ for $v\in\mathcal{V}_{s}$ and generate $\hat{s}$. The parameters of $f_{A}$ are learned using the negative of cross-entropy:
$\mathcal{L}_{A}=$ $\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]$ (3)
$+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))].$
2.3.3 Learning the parameters of BFtS
Let $\theta_{C},\theta_{A}$, and $\theta_{I}$ be the parameters of models $f_{C}$, $f_{A}$, and $f_{I}$, respectively. We propose learning these parameters jointly via a 3-player adversarial learning scheme. Our goal is for $f_{I}$ to learn how to impute the missing sensitive values that represent the worst-fairness scenario for the node representations learned by $f_{C}$, which depend on the sensitive attribute predictions produced by $f_{A}$. This will induce $f_{C}$ to be both fair and accurate. Thus, $\theta_{C}$ are optimized as follows:
$\theta_{C}^{*}=\arg\min_{\theta_{C}}\mathcal{L}_{C}+\alpha\mathcal{L}_{A}$ (4)
where $\alpha$ is a hyperparameter that controls the trade-off between accuracy and fairness of the node classification GNN.
The parameters of the sensitive attribute predictor $\theta_{A}$ only influence the $\mathcal{L_{A}}$ loss. $\theta_{A}$ are learned by simply maximizing the negative of cross-entropy loss:
$\theta_{A}^{*}=\arg\max_{\theta_{A}}\mathcal{L}_{A}$ (5)
Finally, to learn the parameters $\theta_{I}$ of the imputation model toward generating predictions that are accurate and represent the worst-case scenario for the fairness of node representations, we apply the following optimization:
$\theta_{I}^{*}=\arg\min_{\theta_{I}}\mathcal{L}_{I}-\beta\mathcal{L}_{A}$ (6)
Here $\beta$ is a hyperparameter that controls the trade-off between imputation accuracy and the bias generated by the imputed values. The min-max objective between the three players is therefore $\min_{C}\max_{I,A}\mathcal{L}_{A}$.
$\begin{split}\min_{\theta_{C}}\max_{\theta_{I},\theta_{A}}~{}&\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]\\ &+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))].\end{split}$ (7)
See the Supp. Material for a pseudocode for BtFS.
VANILLA DEBIAS FAIRGNN RNF BFtS (Ours)
Bail AVPR ($\uparrow$) 0.84±0.00 0.83±0.00 0.83±0.00 0.81±0.00 0.88±0.00
F1 score ($\uparrow$) 0.88±0.00 0.85 ±0.00 0.86±0.00 0.86±0.01 0.91±0.00
%$\%\Delta DP$ ($\downarrow$) 10.10±0.03 8.78±0.02 7.9±0.03 8.50±0.04 7.60±0.03
%$\%\Delta EQOP$ ($\downarrow$) 5.20±0.04 4.50±0.03 4.10±0.04 4.70±0.05 1.70±0.00
Credit AVPR ($\uparrow$) 0.83±0.00 0.82±0.00 0.81±0.00 0.81±0.00 0.81±0.00
F1 score ($\uparrow$) 0.79±0.00 0.73 ±0.00 0.76±0.00 0.75±0.01 0.76±0.00
%$\Delta DP$ ($\downarrow$) 12.10±0.10 6.80±0.20 5.60±0.08 5.30±0.09 5.40±0.02
%$\%\Delta EQOP$ ($\downarrow$) 14.50±0.10 6.20±0.10 5.90±0.10 5.10±0.00 4.50±0.01
German AVPR ($\uparrow$) 0.75±0.00 0.72 ±0.00 0.73±0.00 0.72±0.01 0.72±0.00
F1 score ($\uparrow$) 0.74±0.01 0.73 ±0.00 0.74±0.00 0.72±0.01 0.74±0.00
%$\Delta DP$ ($\downarrow$) 6.30±0.10 5.02±0.20 5.70 ±0.30 4.20±0.00 2.74±0.02
%$\%\Delta EQOP$ ($\downarrow$) 8.40±0.08 6.24±0.20 5.70±0.20 5.90±0.01 1.70±0.04
NBA AVPR ($\uparrow$) 0.75±0.00 0.73±0.00 0.73±0.00 0.73±0.00 0.74±0.00
F1 score ($\uparrow$) 0.73±0.00 0.67 ±0.00 0.70±0.00 0.69±0.00 0.70±0.00
%$\Delta DP$ ($\downarrow$) 8.10±0.02 3.40±0.04 1.20±0.03 1.50±0.05 1.10±0.04
%$\%\Delta EQOP$ ($\downarrow$) 16.50±0.01 5.50±0.06 2.30±0.00 2.10±0.03 1.70±0.03
Table 1: AVPR, F1 score, %$\%\Delta DP$, and %$\%\Delta EQOP$ of different methods. For the Bail, and NBA dataset, we outperform all baselines in terms of accuracy and fairness. For other datasets, we improve fairness while slightly sacrificing accuracy.
2.4 Theoretical analysis
We will analyze some theoretical properties of BFtS. All the proofs are in the Supp. Material.
Proposition 1.
Given optimal adversaries $f_{A}^{*}$ and $f_{I}^{*}$, the GNN classifier $f_{C}$ solves the following optimization problem:
$\min_{\theta_{C}}\sup_{\hat{s}}JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0)),$ (8)
where $JS$ is the Jensen Shannon divergence.
Theorem 1.
When the GNN classifier $f_{C}$ reaches global minima, it will achieve demographic parity (Eqn 9) for the worst-case imputation (minimum fairness) $\hat{s}$.
We note that different from Dai and Wang (2021), our guarantee is for the worst-case imputation and not the true sensitive value. That is because our imputations $\hat{s}$ are not independent of the representations $\mathbf{h}$ (see Eqn 7).
Corollary 1.
Let $s^{\prime}=f(G,X)$ be an imputation method independent of models $f_{C}$ and $f_{A}$, then $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\leq JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$.
The value of $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is key for the convergence of adversarial learning (see discussion in Supp. Material). For instance, for FairGNN, if $s^{\prime}$ is inaccurate, $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\approx 0$ and $\mathcal{L}_{A}$ in Eqn 4 will be a constant. The corollary shows that our method is more robust to this issue—as also illustrated in our experiments.
3 Experimental Results
Figure 3: Performance of the methods using the Simulation dataset for different values of assortativity coefficients. In the x-axis, we plot the F1 score and in the y axis of the top row we plot $1-DP$ and in the y-axis of the top row we plot $1-EQOP$. The top right corner of the plot, therefore, represents a high F1 with low bias. When the assortativity is low, other methods fail to learn the node labels. With higher assortativity, though other methods learn the class labels, BFtS is less biased for similar accuracy. Note that the X-axes have different ranges.
We compare our approach for learning fair GNNs for node classification (BtFS) against the state-of-the-art in terms of accuracy and fairness using real and synthetic datasets.
Figure 4: Fairness vs. accuracy for all the methods. The x-axis of each plot shows the F1 score. We plot $1-\Delta DP$ and $1-\Delta EQOP$ on the y-axis of rows 1 and 2, respectively. Therefore, the top right corner of the plot represents a high F1 with low bias. Our approach (BFtS) often achieves better fairness for a similar value of F1 compared to the baselines.
Figure 5: 2-D kernel PCA node representations for the Bail dataset generated by the FairGNN and BFtS (our approach). We use different markers for nodes depending on their predicted class, sensitive attribute (race), and whether the sensitive attribute is missing. The results show that representations for nodes with missing values are more biased for FairGNN—they are concentrated in a negative class region. On the other hand, nodes with missing values are better spread over the space for BFtS representations. In BFtS, there are more ”Race = Black” nodes than in FairGNN that have missing sensitive values in training but are predicted to be positive.
3.1 Datasets
The datasets are described in the Supp. Materials.
3.2 Baselines
Vanilla: The vanilla model is the basic GNN model trained with cross-entropy loss without any fairness intervention.
Debias Zhang et al. (2018): Debias is an adversarial learning method that trains a classifier and an adversary simultaneously. The adversary is trained using the softmax output of the last layer of the classifier. It does not handle missing sensitive data. Hence, we only use $\mathcal{V}_{S}\in\mathcal{V}$ to calculate the adversarial loss.
FairGNN Dai and Wang (2021): FairGNN first uses a GNN to estimate the missing sensitive attributes. It then trains another GNN-based classifier and a DNN adversary. The adversary helps to eliminate bias from the representations learned by the GNN classifier.
RNF Du et al. (2021): RNF imputes the missing sensitive attributes by first training the model with generalized cross entropy Zhang and Sabuncu (2018). It eliminates bias from the classification head. RNF uses samples with the same ground-truth label but distinct sensitive attributes and trains the classification head using their neutralized representations.
3.3 Evaluation
Accuracy Evaluation: As with many real-world datasets in fairness literature, our datasets are imbalanced. To evaluate their accuracy, we apply average precision (AVPR) and the F1 score.
Fairness Evaluation: We apply two fairness metrics that will be defined based on a sensitive attribute $s$. Let $s=0$ be the protected group and $s=1$ be the non-protected group. The Demographic Parity (DP) requires the prediction to be independent of the sensitive attribute Kamiran and Calders (2009). We evaluate demographic parity as-
$\Delta DP(\hat{y},s)=|p(\hat{y}=1|s=1)-p(\hat{y}=1|s=0)|$ (9)
Equality of Opportunity (EQOP) calculates the difference between true positive rates between protected and non-protected groups Hardt et al. (2016). We evaluate Equality of Opportunity as-
$\begin{split}\Delta EQOP(\hat{y},s,y)=&|p(\hat{y}=1|s=1,y=1)\\ &-p(\hat{y}=1|s=0,y=1)|.\end{split}$
$\Delta DP$ and $\Delta EQOP$ calculate bias. We can get fairness by subtracting bias from $1$.
3.4 Hyperparameter settings
Each GNN has two layers and a dropout probability of $0.5$. Unless otherwise stated, $\mathcal{V}_{S}$ was set to $40\%$ of training data for all experiments. $\mathcal{V}_{L}$ is $300,400,1400,100$, and $800$ for the Simulation, German, Credit, NBA and Bail datasets, respectively. We adjust the values of $\alpha$ and $\beta$ for each dataset using cross-validation based on average precision. For all baselines, we choose the respective regularizers using cross-validation based on the same metrics. Every model, except RNF, combines the training of several networks with various convergence rates. We stop training RNF based on early stopping to eliminate overfitting. We train the rest of the models for certain epochs and choose the model with the best AVPR. All the results shown in the following section are the average of $10$ runs. We use GCN as the GNN architecture for all experiments unless mentioned otherwise.
3.5 Results and Analysis
Figure 3 shows the F1 score, Demographic Parity, and Equality of Opportunity of different methods while varying their parameter that controls the accuracy vs. fairness trade-off. We have generated the simulation graph using a Stochastic Block Model with different assortativity coefficients—i.e. the extent in which links exist within clusters compared with across clusters. Learning missing sensitive values and node labels with a low assortativity coefficient is hard and graphs with assortativity $0.17$ and $0.37$ represent the scenario described in Corollary 1, therefore, the JS divergence tends to be small. Figure 3 shows that for assortativity $0.17$ and $0.37$, FairGNN and Debias fail to learn the node labels in multiple cases. As we increase the assortativity coefficient, all benchmark models can learn the node labels. BFtS achieves a better fairness vs. accuracy trade-off in all cases.
Table 1 shows the average results for Bail, Credit, German, and NBA datasets. In terms of accuracy and fairness, BFtS outperforms all baseline approaches for the Bail dataset. Bail dataset is highly imbalanced with $7104$ positive samples amongst $18876$. Therefore, LDAM loss of the GNN classifier results in better AVPR and F1 than the Vanilla model. We also outperform competitors on the Credit dataset in terms of fairness with a slight AVPR loss. Compared to all other baselines, our bias is lower for the German dataset. To achieve this fairness, however, we sacrifice some AVPR. Although we lose some AVPR in the German and Credit datasets, we win in F1 and the AVPR loss is negligible when compared to the fairness benefit.
We demonstrate an accuracy and fairness trade-off in Figure 4. For BFtS, we vary $\alpha$, and $\beta$ to control the trade-off between accuracy and fairness. For other models, we vary their respective regularizers. On the x-axis of Figure 4, we have the F1 score. We have $1-\Delta DP$, and $1-\Delta EQOP$ on the y-axis of rows 1 and 2, respectively. Therefore, the top right corner of the plot represents higher fairness and accuracy. Our model achieves better fairness and similar accuracy for the Bail, NBA, German, and Simulation $(assortativity=0.67)$ datasets. For the Credit dataset, BFtS achieves competitive results against RNF and FairGNN.
Figure 5 shows Kernel PCA node representations produced by FairGNN and BFtS using the Bail dataset. We show the samples with missing ”Race” with a navy dot in the middle. The figure shows samples predicted as positive class and negative class in green and red, respectively. The representations generated by FairGNN are noticeably more biased than the ones generated by our approach. We can see a higher number of ”Race = Black” samples that had the sensitive values missing in training are predicted as positive by BFtS than that of FairGNN. BFtS spreads nodes with missing values more uniformly over the space. The missing value imputation of FairGNN underestimated the bias in the dataset and therefore, the model could not overcome the bias in the predictions.
4 Conclusion
We present BFtS, a 3-player adversarial learning framework for the imputation of missing sensitive attributes that produces challenging values for graph-based fairness. Our theoretical analysis showed that our approach is more robust to convergence challenges during adversarial training than existing alternatives. We evaluate the BFtS architecture using synthetic and real datasets. Our experimental results provide strong evidence that our approach achieves a better fairness $\times$ accuracy trade-off than the baselines.
Our work opens several venues for future research. We will investigate methods for estimating the expected fairness based on the uncertainty over missing values. We are also interested in fair graph machine learning under missing links.