### v3

# Better Fair than Sorry: Adversarial Missing Sensitive Value Imputation for Fair GNNs

Abstract
Graph Neural Networks (GNNs) have achieved state-of-the-art results in many relevant tasks where decisions might disproportionately impact specific communities. However, existing work on fair GNNs often assumes that either protected attributes are fully observed or that the missing protected attribute imputation is fair. In practice, biases in the imputation will propagate to the model outcomes, leading them to overestimate the fairness of their predictions. We address this challenge by proposing Better Fair than Sorry (BFtS), a fair missing data imputation model for protected attributes. The key design principle behind BFtS is that imputations should approximate the worst-case scenario for fairness—i.e. when optimizing fairness is the hardest. We implement this idea using a 3-player adversarial scheme where two adversaries collaborate against a GNN-based classifier, and the classifier minimizes the maximum bias. Experiments using synthetic and real datasets show that BFtS often achieves a better fairness $\times$ accuracy trade-off than existing alternatives.
1 Introduction
With the increasing popularity of machine learning models for high-stakes decision-making, it has become a consensus that these models carry implicit biases, and biases should be addressed to improve the fairness of algorithmic decisions Ghallab (2019). The disparate treatment of such models towards African Americans and women has been illustrated in the well-documented COMPAS Angwin et al. (2022) and Apple credit card Vigdor (2019) cases, respectively. While there has been extensive research on fair algorithms and fair machine learning, the proposed solutions have mostly disregarded important challenges that arise in real-world settings.
(a) Missing values.
(b) Imputed sensitive attributes.
(c) Worst-case imputation.
Figure 1: Motivation: In Figure (a), a graph machine learning algorithm is applied to decide who receives credit (positive or negative) based on a possibly missing sensitive attribute, gender (and binary only for illustrative purposes). As shown in Figure (b), traditional missing data imputation does not account for outcomes (positive/negative), and thus, their imputed values can under-represent the bias of the complete dataset—demographic parity (DP) is $0.09$ in this example. This paper proposes BFtS, an imputation method for graph data that optimizes fairness in the worst-case imputation scenario using adversarial learning, as shown in Figure (c), where DP is $0.47$.
Existing work in fair machine learning has mostly focused on tabular, image, and text data. However, in several applications, data can be naturally modeled as graphs (or networks), representing different objects, their relationships, and attributes. Graph Neural Networks (GNNs) have achieved state-of-the-art results in graph machine learning tasks, including node classification, link prediction, and graph classification Gilmer et al. (2017); Kipf and Welling (2017). For instance, in the case of link prediction in a professional network, such as LinkedIn, which is a recruiting and networking tool, link recommendations should not be biased against protected groups Fabbri et al. (2020); Current et al. (2022). However, guaranteeing fairness in graph data is a challenge due to correlations in the connections caused by homophily and influence.
Another challenge for fair machine learning is the pervasiveness of missing data Little and Rubin (2019). This is particularly critical because fair algorithms often require knowledge of sensitive attributes that are more likely to be missing due to biases in the collection process or privacy concerns. As an example, the absence of data in census surveys and health surveys is correlated with sensitive attributes such as gender, age, race, etc. O’Hare and O’Hare (2019); Weber et al. (2021). Missing data is also a concern in networked data, especially when tracking the transmission of diseases Ghani et al. (1998). As a consequence, many fair machine learning algorithms require the use of missing imputation methods to mitigate their biases, and biases in imputed values might be inadvertently propagated to the final decisions Mansoor et al. (2022).
This paper addresses the problem of mitigating bias in graph machine learning in the presence of missing sensitive attributes. We propose an adversarial imputation method that generates challenging instances for a broad class of fair algorithms based on GNNs. Figure 1 motivates our approach. A machine learning algorithm is applied to decide whether individuals should or should not receive credit. Figure 1(a) shows both the gender and outcome for each individual. The genders of some samples are unknown in Figure 1(a), and the demographic parity ($\Delta DP$) considering only the observed values is $0.25$ ($\Delta DP$ is a bias metric described in Section 4.2). We illustrate how different missing data imputations affect the fairness of credit decisions in Figures 1(b) and 1(c). A straightforward imputation is demonstrated in Figure 1(b), where the gender of the majority of neighbors is assigned to the missing attribute, resulting in a best-case scenario in terms of fairness with a $\Delta DP$ of $0.09$. However, in the worst-case scenario, shown in Figure 1(c), $\Delta DP$ is $0.47$. The dataset’s bias might be highly underrepresented in the imputation of Figure 1(b), and the final decisions of a fair algorithm will still be unfair in that case. We show the correlation of sensitive information and true class labels of different datasets (details in supp. material) in Figure 2. We use a Graph Convolutional Network (GCN) to predict the missing sensitive values. When $<50\%$ sensitive information is available, the predicted value underrepresented the true underlying bias in all datasets. Our method, BfTS, can better represent the true bias in the dataset. Our results provide strong evidence that off-the-shelf imputation methods often misguide fair graph machine learning algorithms due to underestimated bias.
Figure 2: Bias in training (i.e., the correlation between the sensitive attribute and node labels) for independent imputation and our approach (BFtS) with varying missing ratios. The rightmost column shows the true bias. Compared with BFtS, independent imputation often underestimates the true bias in the data.
We propose Better Fair than Sorry (BFtS), which is a 3-player adversarial learning framework for missing sensitive value imputation. Figure 3 provides an overview of the proposed approach. The GNN $f_{C}$ interacts with two adversaries $f_{A}$ and $f_{I}$. $f_{A}$ attempts to predict sensitive values based on the representations learned by $f_{C}$ while $f_{I}$ imputes missing values that help $f_{A}$ maximize its accuracy. As a high accuracy of $f_{A}$ indicates high bias in the representations learned by $f_{C}$, $f_{C}$’s parameters are optimized to minimize the bias of its representations in the worst case in terms of fairness. Our experiments show that BFtS achieves a better accuracy $\times$ fairness trade-off than existing approaches. We summarize the contributions of this paper as follows:
  • We propose Better Fair than Sorry (BFtS), a novel 3-player adversarial learning framework for the imputation of missing sensitive data that produces worst-case imputed values for fair GNNs and is effective even when sensitive attribute information is completely unavailable.
  • We prove that the maximal demographic parity (Eq. 8) of BFtS is minimal among all estimators of missing sensitive attributes and that BFtS is robust to convergence issues in adversarial training.
  • We show empirically that BFtS achieves a better fairness $\times$ accuracy trade-off than the existing alternatives.
1.1 Related Work
Fairness in graph machine learning. We divide the work in this area into three main directions (pre-processing, extended objective functions, and adversarial learning). FairOT, FairDrop, FairSIN, EDITS, and Graphair are pre-processing techniques that remove the bias from the graph before training Laclau et al. (2021); Spinelli et al. (2021); Yang et al. (2024); Dong et al. (2022); Ling et al. (2023). Fairwalk, Crosswalk, Debayes, MONET, NIFTY, FairVGNN, PFR-AX, and FairGAE modify the objective function of GNNs to learn fair representations during training Rahman et al. (2019); Khajehnejad et al. (2022); Buyl and De Bie (2020); Palowitch and Perozzi (2019); Agarwal et al. (2021); Wang et al. (2022); Merchant and Castillo (2023); Fan et al. (2021). CFC, FLIP, DKGE, and Debias train adversarial networks along with GNN for fair prediction Bose and Hamilton (2019); Masrour et al. (2020); Arduini et al. (2020); Zhang et al. (2018). However, different from our work, all these methods require fully observed sensitive attributes to ensure fairness.
Missing Data Imputation and fairness. Missing data can be imputed using unconditional mean, reconstruction, preferential attachment, autoencoders, etc. Donders et al. (2006); Pereira et al. (2020); Huisman (2009). The traditional procedure for handling missing data is to first impute the missing values, and then solve the task Rossi et al. (2021); Buck (1960). SAT is a matching-based GNN for graphs with missing attributes Chen et al. (2020), but it does not account for fairness. Training a classifier from imputed data can amplify the unfairness of a machine learning model, as discussed in Subramonian et al. (2022); Zhang and Long (2021); Feng et al. (2023). Some studies try to generate fair feature imputations. Subramonian et al. (2022) has modified mean aggregation feature imputation to minimize discrimination risk against protected groups. Feng et al. (2023) has proposed recursive partitioning and fair uniform resampling for fair classification with missing values. However, as the approaches discussed earlier, these methods require full knowledge of the protected attributes.
Fairness with missing protected attributes. These studies are the ones most related to ours. One extreme assumption is the complete unavailability of sensitive information. Hashimoto et al. (2018) takes into account the worst-case distribution across a given group size, and Lahoti et al. (2020) adversarially reweighs training samples to optimize the average loss. Chai and Wang (2022) employs top $k$ average loss minimization to guarantee fairness. Zhao et al. (2022) minimizes the correlation between the predictions and other features associated with sensitive attributes. Yan et al. (2020) achieves fairness via class re-balancing and Chai et al. (2022) uses soft labels from an overfitted teacher model. In practice, some sensitive information is often available and can significantly improve fairness Chai and Wang (2022). FairGNN Dai and Wang (2021) assumes limited sensitive information and learns a fair GNN by predicting the missing values with an independent imputation model. RNF Du et al. (2021) generates proxy annotations for missing attributes using generalized cross-entropy Zhang and Sabuncu (2018). Similar to RNF, our approach (BFtS) handles completely or partially missing sensitive information. BFtS applies a 3-player scheme to minimize the maximum possible bias and outperforms FairGNN and RNF in terms of fairness $\times$ accuracy. Minimizing the maximum instead of the average risk has been shown to achieve better guarantees Shalev-Shwartz and Wexler (2016). A similar minmax approach has been applied to maximize the robustness and accuracy of uncertainty models Löfberg (2003); Lanckriet et al. (2002); Chen et al. (2017); Fauß et al. (2021). In Nguyen et al. (2017); Vandenhende et al. (2019), a 3-player adversarial network is proposed to improve the classification and stability of adversarial learning.
2 Problem Definition and Preliminaries
In this section, we formalize the problem addressed in this paper and introduce the Graph Neural Network model, which is our base model for node classification. For simplicity, we only consider binary-sensitive attributes. By modifying the loss function in Eq. 2 to categorical cross entropy and MSE, BFtS can be generalized for non-binary and continuous protected attributes, respectively.
2.1 Notation and Problem Definition
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X},\mathcal{S})$ be an undirected graph where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, $\mathcal{X}$ are node attributes, and $\mathcal{S}$ is the sensitive attribute. The matrix $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of $\mathcal{G}$ where $A_{uv}=1$ if there is an edge between $u$ and $v$ and $A_{uv}=0$, otherwise. We focus on the setting where sensitive attributes might be missing and only $\mathcal{V}_{S}\subseteq\mathcal{V}$ nodes include the information of the sensitive attribute $s_{v}$ for a node $v$. The sensitive attribute forms two groups, which are often called sensitive ($s_{v}=1$) and non-sensitive ($s_{v}=0$).
We consider the fair node classification task, where the goal is to learn a classifier $f_{C}$ to predict node labels $y_{v}\in\{0,1\}$ based on a training set $\mathcal{V}_{L}\subseteq\mathcal{V}$. Without loss of generality, we assume that the class $y=1$ is the desired one (e.g., receive credit or bail). Given a classification loss $\mathcal{L}_{C}$ and a fairness loss $\mathcal{L}_{A}$, the goal is to minimize their combination:
$\min_{f_{C}}\mathcal{L}_{C}(f_{C})+\alpha\mathcal{L}_{A}(f_{C})$
Computing fairness loss $\mathcal{L}_{A}$ requires sensitive attributes $s_{v}$. The main challenge addressed in this paper is due to our assumption that sensitive attributes can be missing. In this scenario, one can apply an independent imputation model $s_{v}\approx f_{I}(X[v])$ to predict missing sensitive attributes before training the fair classifier. However, the fairness of the resulting classifier will be highly dependent on the accuracy of $f_{I}$, which is the key motivation for our work. In particular, we consider two scenarios to motivate our solution:
Sensitive attributes missing completely at random (MCR) at a high rate:
The missingness process can be modeled by a Bernoulli distribution $Bern(\beta)$ with $\beta$ close to $1$ and thus $f_{I}$ is expected to have a high error rate. Thus, the bias $\mathcal{L}_{A}$ can be underestimated by the fair classification model.
Adversarily missing sensitive attributes:
The missingness process is generated by an adversary that manipulates $f_{I}$ to artificially decrease the fairness loss $\mathcal{L}_{A}$. As an example, the adversary might select missing values to flip the predictions of $s_{v}$ for nodes that are both in the protected group and the undesired class ($y=0$). These scenarios are particularly relevant in graphs as they can exploit structural graph properties (see Sec. 4.1).
As an independent imputation cannot be trusted in these scenarios, we propose an interactive (3-player) worst-case imputation model for fair classification described in Sec. 3.
2.2 Graph Neural Network
Two of the models (or players) learned by our model are GNNs. Without loss of generality, we will assume that they are message-passing neural networks Gilmer et al. (2017), which we will simply refer to as GNNs. A GNN learns node representations $\mathbf{h}_{v}$ for each node $v$ that can be used to predict (categorical or continuous) values $y_{v}=f(\mathbf{h}_{v})$. These representations are learned via successive $\mathsf{AGGREGATE}$ and $\mathsf{COMBINE}$ operations defined as follows:
$\mathbf{h}_{v}^{k}\!=\!\mathsf{COMBINE}^{k}(\mathbf{h}_{v}^{k-1},\mathsf{AGGREGATE}^{k-1}({\mathbf{h}_{u}^{k-1},A_{u,v}\!=\!1})).$
where $\mathbf{a}_{v}^{k}$ aggregates the messages received by the neighbors of $v$ in $\mathcal{G}$ and $\mathbf{h}_{v}^{k}$ are representations learned at the $k$-th layer.
Figure 3: Proposed 3-player framework for fair GNN training with missing data imputation (BFtS). $f_{C}$ generates node representations by minimizing the classification loss $\mathcal{L}_{C}$ (Eqn. 1) and maximizing sensitive attribute prediction loss $\mathcal{L}_{A}$ (Eqn. 3). $f_{A}$ predicts sensitive attributes using the representations from $f_{C}$ by minimizing $\mathcal{L}_{A}$. $f_{I}$ predicts missing values by minimizing the imputation loss $\mathcal{L}_{I}$ (Eqn. 2) and maximizing $\mathcal{L}_{A}$. $\hat{y}$, $\hat{si}$, $\hat{sa}$ are the predictions generated by $f_{C}$, $f_{I}$ and $f_{A}$, respectively.
3 Fairness-Aware Adversarial Missing Data Imputation for Graphs
We introduce a graph learning framework, which we call BFtS (Better Fair than Sorry), that allows fair GNN training in a setting with missing sensitive data. BFtS applies a 3-player adversarial learning scheme, where the fair GNN is trained jointly with two adversaries—one to predict sensitive attributes based on GNN embeddings and another to impute missing values that minimize fairness. As a result, fairness assessments by the GNN are based on the worst-case imputed values generated by the adversarial imputation.
3.1 Proposed Model (BFtS)
Our goal is to learn a fair and accurate GNN $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ for node classification with missing sensitive values. The proposed solution is model-agnostic and based on adversarial learning. It formulates the imputation model as a second adversary of the GNN.
Figure 3 depicts the flow diagram of the proposed model. The model has three primary components: a missing sensitive attribute imputation GNN $f_{I}$, a node classification GNN $f_{C}$, and a sensitive attribute Deep Neural Net (DNN) prediction $f_{A}$. $f_{C}$ takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and predicts the node labels. $f_{A}$ is an adversarial neural network that attempts to estimate the sensitive information from the final layer representations of $f_{C}$ to assess the bias. More specifically, it $f_{C}$ is assumed to be biased if the adversary $f_{A}$ can accurately predict the sensitive attribute information from the representations of $f_{C}$. To train $f_{C}$, we need to know the values of the missing sensitive attributes. $f_{I}$ predicts the missing sensitive attributes. It takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and generates the missing sensitive attributes $\hat{si}$. $f_{I}$ aims to generate sensitive values that minimize the fairness of $f_{C}$. Therefore, it works as a second adversary to the model.
We will formalize the architecture of the players, define their loss functions, and describe the training procedure for our approach in the next section.
3.1.1 Player architectures
This section describes the architecture of each player in BFtS.
GNN classifier $f_{C}$:
Node classification model $\hat{y}_{v}=f_{C}(x_{v},\mathcal{G})$ implemented using a GNN. We assume that $f_{C}$ does not apply sensitive attributes but uses other attributes in $\mathcal{X}$ that might be correlated with sensitive ones. The goal of $f_{C}$ is to achieve both accuracy and fairness. To improve fairness, $f_{C}$ tries to minimize the loss of adversary $f_{A}$.
Sensitive attribute predictor (Adversary 1) $f_{A}$:
Neural network that applies representations $\mathbf{h}_{v}$ generated by $f_{C}$ to predict sensitive attributes as $\hat{sa}_{v}=f_{A}(\mathbf{h}_{v})$. $f_{C}$ is fair if $f_{A}$ performs poorly.
Missing data imputation GNN (Adversary 2) $f_{I}$:
Predicts missing sensitive attributes as $\hat{si}_{v}=f_{I}(x_{v},\mathcal{G})$. However, besides being accurate, $f_{I}$ plays the role of an adversary to $f_{C}$ by predicting values that maximize the accuracy of $f_{A}$.
3.1.2 Loss Functions
Here, we describe the loss functions applied in our approach.
Node Classification Loss
Let $\hat{y}=f_{C}(X,G)$ be the predictions of $f_{C}$. We apply the cross-entropy loss to $f_{C}$’s predictions as follows:
$\mathcal{L}_{C}=-\frac{1}{|\mathcal{V_{L}}|}\sum_{v\in\mathcal{V_{L}}}\hat{y}_{v}log(\hat{y}_{v})+(1-\hat{y}_{v})log(1-\hat{y}_{v}).$ (1)
Sensitive Attribute Imputation Loss
Because sensitive attributes tend to be imbalanced in real datasets, $f_{I}$ applies the Label-Distribution-Aware Margin (LDAM) loss Cao et al. (2019) for classification. Let $\hat{si}=f_{I}(X,G)$ be the one-hot encoded predictions and $\hat{si}^{s}_{v}$ be the output of node $v$ for $s_{v}$. Then, the LDAM loss of $f_{I}$ is defined as:
$\mathcal{L}_{I}=\frac{1}{|\mathcal{V_{S}}|}\sum_{v=1}^{|\mathcal{V}_{S}|}-log\frac{e^{\hat{si}^{s}_{v}-\Delta^{s}}}{e^{\hat{si}^{s}_{v}-\Delta^{s}}+\sum_{k\neq s}e^{\hat{si}^{k}_{v}}}.$ (2)
where $\Delta^{j}=C/n_{j}^{\frac{1}{4}}$ and $j\in\{0,1\}$.
Here $C$ is a constant independent of the sensitive attribute $s$, and $n_{j}$ is the number of samples that belong to $s=j$. The LDAM loss is a weighted version of the negative log-likelihood loss. Intuitively, since $\Delta^{j}$ is larger for smaller values of $n_{j}$, it ensures a higher margin for the smaller classes.
Furthermore, BFtS can generate fair outputs when very little or no sensitive information is provided. When no or very little sensitive information is available, ($|\mathcal{V}_{S}|\to 0$), $f_{I}$ can impute the sensitive values from the ground truth label $y$ with LDAM loss. We can replace $\mathcal{V}_{S}$ with $\mathcal{V}_{L}$ and $s$ with $y$ for each node $v\in\mathcal{V}_{L}$ in Eqtn 2. The primary argument is that the worst-case fairness model generally assigns a more desired outcome to the non-protected group and a less desired outcome to the protected group. Therefore, the nodes predicted as the minority class will fall into the protected group, and vice versa. Compared to the model $f_{I}$ trained with LDAM loss with limited sensitive information of $\mathcal{V}_{S}$; without any sensitive information, $f_{I}$ trained with LDAM loss learns more bias-amplified samples with labels $y$ of $V_{L}$. Depending on the amount of sensitive information available, the user can choose any of these two methods to learn the imputation.
Sensitive Attribute Prediction Loss
Given the sensitive information for $\mathcal{V}_{s}$, we first replace $\hat{si}_{v}$ by ${s}_{v}$ for $v\in\mathcal{V}_{s}$ and thereby generate $\hat{s}$. The parameters of $f_{A}$ are learned using:
$\mathcal{L}_{A}=$ $\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]$
$+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))]$ (3)
3.1.3 Learning the parameters of BFtS
Let $\theta_{C},\theta_{A}$, and $\theta_{I}$ be the parameters of models $f_{C}$, $f_{A}$, and $f_{I}$, respectively. We propose learning these parameters jointly via a 3-player adversarial learning scheme. Our goal is for $f_{I}$ to learn how to impute the missing sensitive values that represent the worst-fairness scenario for the node representations learned by $f_{C}$, which depend on the sensitive attribute predictions produced by $f_{A}$. This will induce $f_{C}$ to be both fair and accurate. Thus, $\theta_{C}$ are optimized as follows:
$\theta_{C}^{*}=\arg\min_{\theta_{C}}\mathcal{L}_{C}+\alpha\mathcal{L}_{A}.$ (4)
where $\alpha$ is a hyperparameter that controls the trade-off between accuracy and fairness.
The parameters of the sensitive attribute predictor $\theta_{A}$ only influence the $\mathcal{L_{A}}$ loss. $\theta_{A}$ are learned by maximizing $\mathcal{L}_{A}$:
$\theta_{A}^{*}=\arg\max_{\theta_{A}}\mathcal{L}_{A}.$ (5)
To learn the parameters $\theta_{I}$ of the imputation model to generate predictions that are accurate and represent the worst-case scenario for fairness, we apply the following:
$\theta_{I}^{*}=\arg\min_{\theta_{I}}\mathcal{L}_{I}-\beta\mathcal{L}_{A}.$ (6)
Here $\beta$ is a hyperparameter that controls the trade-off between imputation accuracy and worst-case imputation. The min-max objective between the three players is therefore:
$\begin{split}\min_{\theta_{C}}\max_{\theta_{I},\theta_{A}}~{}&\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]\\ &+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))].\end{split}$ (7)
3.2 Training Algorithm
Algorithm 1 is a high-level description of the key steps applied for training BFtS. It receives the graph $\mathcal{G}$, labels $y$, labeled nodes $\mathcal{V}_{L}$, nodes with observed protected attributes $\mathcal{V}_{S}$ and hyperparameters, $\alpha$ and $\beta$ as inputs and outputs the GNN classifier $f_{C}$, sensitive attribute predictor, $f_{A}$ and missing data imputation GNN $f_{I}$. We first estimate $\hat{si}$ and then fix $\theta_{C}$ and $\theta_{A}$ to update $\theta_{I}$. Then we update $\theta_{A}$ with Eq. 5. After that, we fix $\theta_{A}$ and $\theta_{I}$ and update $\theta_{C}$ with Eq. 4 and repeat these steps until convergence. The time complexity analysis of the algorithm is discussed in the Supp. material.
Algorithm 1 Training BFtS
Input: $\mathcal{G}$; $y$; $\mathcal{V}_{L}$; $\mathcal{V}_{S}$; $\alpha$, $\beta$
Output: $f_{C}$, $f_{A}$, $f_{I}$
1:repeat
2:     Get the estimated sensitive attributes $\hat{si}$ with $f_{I}$
3:     Fix $\theta_{C}$ and $\theta_{A}$ and update $\theta_{I}$ using Eq. 6
4:     Update $\theta_{A}$ using Eq. 5
5:     Fix $\theta_{I}$ and $\theta_{A}$ and update $\theta_{C}$ using Eq. 4
6:until converge
3.3 Theoretical analysis
We will analyze some theoretical properties of BFtS. All the proofs are provided in the Supp. material.
Theorem 1.
The optimal $f^{*}_{I}$ generates the worst-case (minimum fairness) imputation so that $\Delta DP_{f_{C},f^{*}_{I}}\geq\Delta DP_{f_{C},f_{I}}$ where, $\Delta DP_{f_{C},f_{I}}=|p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|\hat{s}=0)|$, $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ and, $\hat{s}=f_{I}(\mathcal{G},\mathcal{X}).$
$f_{I}$ generates imputations such that $\Delta DP$ in Eqn 8 is maximized for a given classifier $f_{C}$
Theorem 2.
$f_{C}$ is a minimax estimator, its maximal $\Delta DP$ is minimal among all estimators of $s$.
The optimal GNN classifier $f^{*}_{C}$, will achieve demographic parity ($\Delta DP=0$ in Eqn 8) for the worst-case imputation (minimum fairness) $\hat{s}$ generated by $f_{I}$.
Corollary 2.1.
Let $s^{\prime}=f(G,X)$ be an imputation method independent of models $f_{C}$ and $f_{A}$, then the Jensen Shannon divergence $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\leq JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$.
The value of $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is related to the convergence of adversarial learning. For FairGNN, if $s^{\prime}$ is inaccurate, $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\approx 0$ and $\mathcal{L}_{A}$ in Eq. 4 will be constant. The interplay between the three players makes BFtS more robust to convergence issues because the objective minimizes the upper bound on the JS divergence. This reduces the probability that the divergence vanishes during training (see details in Supp. material).
4 Experiments and Results
Figure 4: Performance of the methods using the Simulation dataset for different values of assortativity coefficients. In the x-axis, we plot the F1 score, and in the y-axis of the top row, we plot $1-\Delta DP$, and in the y-axis of the bottom row, we plot $1-\Delta EQOP$. The top right corner of the plot, therefore, represents a high F1 with low bias. When the assortativity is low, other methods fail to learn the node labels. With higher assortativity, though other methods learn the class labels, BFtS is less biased and has similar accuracy. Note that the X-axes have different ranges.
We compare our approach for learning fair GNNs for node classification (BtFS) against the state-of-the-art in terms of accuracy and fairness using real and synthetic data. Our code can be found in an anonymous repository [^1] https://anonymous.4open.science/r/BFtS-6ADA. The details of the datasets, baselines, and hyperparameters are in the Supp. material. Results with graph attention network and ablation study are also in the Supp. material.
Figure 5: Fairness vs. accuracy for all the methods. The x-axis of each plot shows the F1 score. We plot $1-\Delta DP$ and $1-\Delta EQOP$ on the y-axis of rows 1 and 2, respectively. Therefore, the top right corner of the plot represents a high F1 with low bias. Our approach (BFtS) often achieves better fairness for a similar value of F1 compared to the baselines.
RNF BFtS
AVPR($\uparrow$) F1 Score($\uparrow$) %$\%\Delta DP$($\downarrow$) %$\%\Delta EQOP(\downarrow$) AVPR($\uparrow$) F1 Score($\uparrow$) %$\%\Delta DP$ ($\downarrow$) %$\%\Delta EQOP(\downarrow)$
Bail 0.81±0.00 0.85±0.00 11.65±0.02 8.51±0.02 0.83±0.00 0.86±0.00 8.01±0.03 4.12±0.01
Credit 0.80±0.00 0.76±0.00 8.01±0.08 7.25±0.08 0.82±0.00 0.77±0.00 5.97±0.10 4.96±0.06
German 0.73±0.00 0.74±0.00 9.42±0.04 8.92±0.12 0.74±0.00 0.75±0.00 6.42±0.11 7.68±0.17
NBA 0.70±0.00 0.70±0.00 6.47±0.03 5.89±0.05 0.73±0.00 0.70±0.00 5.19±0.01 3.59±0.05
Pokec-z 0.73±0.00 0.71±0.00 6.18±0.05 6.29±0.07 0.73±0.00 0.73±0.00 5.10±0.02 3.20±0.06
Pokec-n 0.72±0.00 0.72±0.00 7.58±0.01 7.09±0.04 0.73±0.00 0.70±0.00 4.29±0.02 3.01±0.03
Table 1: AVPR, F1 score, %$\%\Delta DP$, and %$\%\Delta EQOP$ without any protected information of BFtS and RNF. Amongst the baselines, only RNF works without any protected information. BFtS outperforms RNF in terms of fairness and accuracy for all baselines.
4.1 Selection of Missing Value
We generate missing values synthetically in our experiments. To simulate an adversarial missingness process, we select low-degree nodes to have missing values. This choice is based on previous work showing that low-degree nodes have lower prediction accuracy Tang et al. (2020); Liu et al. (2023); Ju et al. (2024). To generate $k$ nodes with missing sensitive data, we choose the $k$ nodes with the lowest degree. The size of $\mathcal{V}_{S}$ was set to $30\%$ of $|\mathcal{V}|$ unless stated otherwise.
4.2 Evaluation
Accuracy Evaluation: As with many datasets in the fairness literature, our datasets are imbalanced. To evaluate their accuracy, we apply average precision (AVPR) and the F1 score.
Fairness Evaluation: We apply two bias metrics that will be defined based on a sensitive attribute $s$. Let $s=0$ be the protected group and $s=1$ be the non-protected group. The Demographic Parity (DP) requires the prediction to be independent of the sensitive attribute Kamiran and Calders (2009). We evaluate demographic parity as:
$\Delta DP(\hat{y},s)=|p(\hat{y}=1|s=1)-p(\hat{y}=1|s=0)|.$ (8)
Equality of Opportunity (EQOP) calculates the difference between true positive rates between protected and non-protected groups Hardt et al. (2016) as follows:
$\begin{split}\Delta EQOP(\hat{y},s,y)=&|p(\hat{y}=1|s=1,y=1)\\ &-p(\hat{y}=1|s=0,y=1)|.\end{split}$ (9)
4.3 Results and Analysis
Figure 4 shows the F1 score, $1-\Delta DP$, and $1-\Delta EQOP$ for different methods while varying their hyperparameter. We have generated the simulation graph using a stochastic block model with different assortativity coefficients, i.e., the extent to which links exist within clusters compared with across clusters. Learning missing sensitive values and node labels with a low assortativity coefficient is hard and graphs with assortativity $0.17$ and $0.37$ represent the scenario described in Corollary 1, therefore, the JS divergence tends to be small for independent imputation, which may result in a lack of convergence for the baselines. Figure 4 shows that for assortativity $0.17$ and $0.37$, FairGNN and Debias fail to predict node labels in multiple cases. As we increase the assortativity, all baselines can predict the labels, but BFtS still achieves a better fairness vs. accuracy trade-off in all cases.
Figure 6: 2-D kernel PCA node representations for the Bail dataset generated by the FairGNN and BFtS (our approach). We use different markers for nodes depending on their predicted class, sensitive attribute (race), and whether the sensitive attribute is missing. The results show that representations for nodes with missing values are more biased for FairGNN where they are concentrated in a negative class region. On the other hand, nodes with missing values are better spread over the space for BFtS representations. In BFtS, there are more ”Race = Black” nodes than in FairGNN that have missing sensitive values in training but are predicted to be positive.
We demonstrate the accuracy vs. fairness trade-off for real datasets in Figure 5. For BFtS, we vary $\alpha$ and $\beta$ to generate different values of fairness and accuracy. For other models, we vary their respective regularizers. The top right corner of the plot represents high fairness and accuracy. Our model achieves better fairness and similar accuracy to the best baseline for the Bail, NBA, German, pokec-n and pokec-z. For all datasets, BFtS achieves a better fairness-accuracy tradeoff.
Figure 6 shows kernel PCA node representations produced by FairGNN and BFtS using the Bail dataset. We show the samples with missing ”Race” with a navy dot in the middle and samples predicted as positive class and negative class in green and red, respectively. The representations generated by FairGNN are noticeably more biased than the ones generated by our approach. A higher number of ”Race = Black” samples with missing sensitive values in training are predicted as positive by BFtS than for FairGNN. Moreover, BFtS spreads nodes with missing values more uniformly over the space. This illustrates how the missing value imputation of FairGNN underestimates the bias in the training data and, therefore, the model could not overcome the bias in the predictions.
Table 1 shows the accuracy and fairness results for Bail, Credit, German, Pokec-z, Pokek-n and NBA without any sensitive attribute information. Amongst the baselines, only RNF works in this setting. The results show that BFtS outperforms RNF for the Bail, German, and Credit. Our approach also outperforms RNF in terms of fairness on Pokec-z,Pokec-n, and NBA with similar AVPR and F1. As expected, we note that when we have $30\%$ of the sensitive attributes available (Figure 5), we obtain higher fairness than when we have no sensitive attribute information (Table 1).
Table 2 shows the imputation accuracy of the methods. We exclude results for Debias, which is based only on the sensitive information available, and for FairVGNN and FairSIN, as they are identical to FairGNN (same imputation method). The results show that BFtS imputation outperforms independent imputation methods. BFtS worst-case imputation based on the LDAM loss outperforms the alternatives in terms of accuracy. This is due to the adversarial missingness process where low-degree nodes are selected to have missing values. Independent imputation methods are less effective in this adversarial setting.
German Credit Bail NBA pokec-z pokec-n
FairGNN 0.70 0.90 0.56 0.78 0.81 0.84
RNF 0.59 0.73 0.53 0.64 0.65 0.68
BFtS 0.72 0.94 0.63 0.78 0.83 0.87
Table 2: Accuracy of missing sensitive imputation.
5 Conclusion
We present BFtS, a 3-player adversarial learning framework for the imputation of missing sensitive attributes that produce challenging values for graph-based fairness. Our theoretical analysis showed that our approach is more robust to convergence challenges during adversarial training than existing alternatives. We evaluate BFtS using synthetic and real datasets. Our experimental results provide strong evidence that our approach achieves a better fairness $\times$ accuracy trade-off than the baselines.
Our work opens several venues for future research. We will investigate methods for estimating the expected fairness based on the uncertainty over missing values. We are also interested in fair graph machine learning under missing links.